# Long-Tailed 3D Detection via 2D Late Fusion

### v3

@@ Abstract
Long-Tailed 3D Object Detection (LT3D) addresses the problem of accurately detecting objects from both common and rare classes. Contemporary multi-modal detectors achieve low AP on rare-classes (e.g., CMT only achieves 9.4 AP on stroller), presumably because training detectors end-to-end with significant class imbalance is challenging. To address this limitation, we delve into a simple late-fusion framework that ensembles independently trained uni-modal LiDAR and RGB detectors. Importantly, such a late-fusion framework allows us to leverage large-scale uni-modal datasets (with more examples for rare classes) to train better uni-modal RGB detectors, unlike prevailing multimodal detectors that require paired multi-modal training data. Notably, our approach significantly improves rare-class detection by 7.2% over prior work. Further, we examine three critical components of our simple late-fusion approach from first principles and investigate whether to train 2D or 3D RGB detectors, whether to match RGB and LiDAR detections in 3D or the projected 2D image plane for fusion, and how to fuse matched detections. Extensive experiments reveal that 2D RGB detectors achieve better recognition accuracy for rare classes than 3D RGB detectors and matching on the 2D image plane mitigates depth estimation errors. Our late-fusion approach achieves 51.4 mAP on the established nuScenes LT3D benchmark, improving over prior work by 5.9 mAP!
Keywords: Long-Tailed 3D Detection, Multimodal Fusion, 2D Late Fusion
@@ 1 Introduction
3D object detection is an integral component of the autonomous vehicle (AV) perception stack [1, 2, 3, 4]. AVs must accurately detect objects from both common and rare classes for safe navigation. Although contemporary end-to-end trained multi-modal detectors have shown impressive performance on common classes (e.g., car and bus), they still struggle to detect rare classes (e.g., stroller and debris) due to limited paired training data in-the-tail. In this paper, we explore the established research problem of Long-Tailed 3D Detection (LT3D) [5], which focuses on achieving high accuracy on both common and rare classes.
@@
Figure 1: We explore a simple late-fusion framework for LT3D by ensembling RGB and LiDAR uni-modal detectors [5]. We rigorously examine three critical components (Fig. 2) from first principles and propose a simple method that fuses detections produced by 2D RGB and 3D LiDAR detectors. Our method achieves 51.4 mAP on the nuScenes [2] LT3D benchmark, significantly outperforming prior work [6] by 5.9 mAP (cf. Table 1).
Status quo. LT3D is not well addressed by simply training state-of-the-art multi-modal detectors on both common and rare classes [5]. For example, BEVFusion [6], an end-to-end trained multi-modal transformer-based detector, only achieves 4.4 AP on the rare class child (cf. Table 7). In contrast, Peri et al. [5] finds that late-fusion (cf. Fig. 1) of monocular 3D RGB detections [7] and 3D LiDAR detections [8] improves rare class recognition, achieving state-of-the-art performance on the nuScenes [2] LT3D benchmark. Importantly, [5] demonstrates that ensembling uni-modal detectors is better suited for LT3D because 3D LiDAR detectors achieve high recall but struggle to correctly recognize rare classes, while RGB detectors excel at recognition but are unable to reliably estimate depth. This motivates our work, which investigates late-fusion from first principles.
The devil is in the details. To address LT3D, we delve into the simple late-fusion framework proposed by [5] (cf. Fig. 1) and study three critical design choices from first principles (cf. Fig. 2), including whether to train 2D or 3D RGB detectors, whether to match RGB and LiDAR detections in the 2D image plane or 3D bird’s-eye-view, and how to fuse matched detections.
First, we evaluate the impact of using 2D versus 3D RGB detectors for late-fusion, and find that the former is straightforward to train (without requiring expensive 3D cuboid annotations on RGB), can readily leverage external datasets that have 2D annotations (particularly improving rare class accuracy), and leads to significantly better LT3D performance (cf. Table 2). This is practically meaningful because annotating 2D boxes on RGB images is significantly cheaper than annotating 3D cuboids.
Next, we study the impact of matching 2D RGB detections and 3D LiDAR detections on the 2D image plane versus in the 3D bird’s-eye-view (BEV). Matching detections in the 3D BEV requires inflating 2D detections to 3D using depth imputed from LiDAR points [9], introducing additional depth estimation errors. Instead, we demonstrate that projecting 3D LiDAR detections to the 2D image plane for matching is more robust (cf. Table 2).
Lastly, we explore score calibration and Bayesian fusion to combine matched detections. We find that proper score calibration improves rare class detection and enables probabilistic fusion [10] of LiDAR and RGB detections (cf. Table 3). Without score calibration, rare class detections are often suppressed by overlapping common class detections.
Contributions. We extensively study late-fusion from first principles, identify best practices including using 2D RGB detectors for fusion, and derive a simple method that incorporates the above insights to resoundingly outperform prior works on the established nuScenes and Argoverse 2 LT3D benchmarks. We rigorously ablate each component and justify our design choices. Importantly, the novelty of our work lies in our rigorous exploration of late-fusion in the context of long-tailed 3D detection.
@@
Figure 2: We examine three key components in the late-fusion of uni-modal RGB and LiDAR detectors from first principles: A. whether to train 2D or 3D RGB detectors, B. whether to match uni-modal detections on the 2D image plane or in the 3D bird’s-eye-view (BEV), and C. how to best fuse matched detections. Perhaps surprisingly, our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining calibrated scores with Bayesian fusion yields state-of-the-art LT3D performance, significantly outperforming end-to-end trained multi-modal detectors (cf. Table 1).
@@ 2 Related Work
3D detection for autonomous vehicles (AVs) can be broadly classified based on input modalities: LiDAR-only, RGB-only, and multi-modal detectors. Recent work in 3D detection is inspired by prior work in 2D detection [11, 12, 13]. LiDAR-based detectors like PointPillars [14], CBGS [15], and PVRCNN++ [16] adopt an SSD-like architecture [12] that regresses amodal bounding boxes from a BEV feature map. More recently, CenterPoint [8] adopts a center-regression loss inspired by [11]. Despite significant progress, LiDAR-based detectors often produce many false positives because it is difficult to distinguish foreground objects from background with sparse LiDAR points alone.
@@
On the other hand, monocular RGB-based methods have gained increased interest in recent years due to low sensor cost and wide-spread adoption [17]. FCOS3D [7] extends FCOS [18] by additionally regressing the size, depth, and rotation for each detection. Recent methods such as BEVDet [19, 20] and BEVFormer [21] construct a BEV feature-map by estimating per-pixel depth [22]. PolarFormer [23] introduces a polar-coordinate transformation that improves near-field detection. Despite recent advances in LiDAR-only and RGB-only 3D detectors, multi-modal fusion is essential for LT3D [5] (detailed next). Importantly, using both RGB (for better recognition) and LiDAR (for better 3D localization) helps detect both rare and common classes. We delve into the late-fusion framework described in Fig. 1 to determine how to effectively fuse RGB and LiDAR uni-modal detectors for LT3D (cf. Fig. 2).
@@
Multimodal fusion for 3D object detection is an active research area. Popular approaches can be categorized into input-fusion, feature-fusion, and late-fusion. Input-fusion methods typically augment LiDAR points using image-level features. For example, PointPainting [24] projects LiDAR points onto a semantic segmentation mask and appends corresponding class scores to each point. MVP [25] densifies regions of LiDAR sweeps that correspond with objects in semantic segmentation masks. Frustum PointNets [26] and SRDL [27] leverage 2D RGB detections to regress 3D bounding boxes for LiDAR points within the 2D detection frustum using PointNets [28].
@@
Recent works show that feature-fusion can be more effective than input-fusion. PointFusion [29] fuses global image and point-cloud features prior to detection and MSMDFusion [30] fuses LiDAR and RGB features at multiple scales. TransFusion [31] and BEVFusion [6] fuse features in the BEV space using multi-head attention. Despite the success of transformers for detecting common objects, [5] finds that TransFusion struggles to detect rare classes, as the transformer architecture (adopted in TransFusion and BEVFusion) suffers from limited paired RGB-LiDAR training data in the long tail. For transformers to work well in practice, they may require diverse, large-scale training datasets [32, 33, 34]. Therefore, we opt to study late-fusion of uni-modal detectors, which does not require paired RGB-LiDAR training data.
CLOCs [35] is a late-fusion method that learns a standalone fusion network to merge RGB and LiDAR detections, showing promising results for 3D detection. More recently, [5] introduces a simple non-learned filtering algorithm that effectively removes false-positive LiDAR-detections that are far away from any 3D RGB detections. We delve into this simple late-fusion framework and present a method that significantly outperforms prior methods for LT3D.
Long-tailed detection is not unique to the AV domain and has been well studied in 2D [36]. Existing methods propose reweighting losses [37, 38, 39, 40, 41], rebalancing data sampling [42, 43, 44, 15], balancing gradients computed from imbalanced classes [45], and balancing network weights [46]. CBGS [15] explicitly addresses rare-class 3D detection by up-sampling LiDAR sweeps with instances of rare classes, and pasting rare objects copied from different scenes. More recently, [47] adopts data augmentation via sampling, and [48] uses active learning and hard example mining to obtain more data for rare classes. LT3D presents distinct opportunities and challenges compared to 2D long-tailed detection because LiDAR sensors offer direct geometric and ego-motion cues that are difficult to extract from 2D images.
3 Late-Fusion of RGB and LiDAR Detections
We first analyze different ways of incorporating RGB information in Sec. 3.1, present a simple algorithm for matching RGB and LiDAR detections in Sec. 3.2, and finally describe score calibration and probabilistic fusion in Sec. 3.3.
3.1 How Do We Incorporate RGB Information?

### v4

# Long-Tailed 3D Detection via Multi-Modal Late-Fusion

@@ Abstract
Contemporary autonomous vehicle (AV) benchmarks have advanced techniques for training 3D detectors, particularly on large-scale multi-modal (LiDAR + RGB) data. Surprisingly, although semantic class labels naturally follow a long-tailed distribution, existing benchmarks only focus on a few common classes (e.g., pedestrian and car) and neglect many rare but crucial classes (e.g., emergency vehicle and stroller). However, AVs must reliably detect both common and rare classes for safe operation in the open world. We address this challenge by formally studying the problem of Long-Tailed 3D Detection (LT3D), which evaluates all annotated classes, including those in-the-tail. We address LT3D with hierarchical losses that promote feature sharing across classes, and introduce diagnostic metrics that award partial credit to “reasonable” mistakes with respect to the semantic hierarchy (e.g., mistaking a child for an adult). Further, we point out that rare-class accuracy is particularly improved via multi-modal late fusion (MMLF) of independently trained uni-modal LiDAR and RGB detectors. Importantly, such an MMLF framework allows us to leverage large-scale uni-modal datasets (with more examples for rare classes) to train better uni-modal detectors, unlike prevailing end-to-end trained multi-modal detectors that require paired multi-modal data. Finally, we examine three critical components of our simple MMLF approach from first principles and investigate whether to train 2D or 3D RGB detectors for fusion, whether to match RGB and LiDAR detections in 3D or the projected 2D image plane, and how to fuse matched detections. Extensive experiments reveal that 2D RGB detectors achieve better recognition accuracy for rare classes than 3D RGB detectors and matching on the 2D image plane mitigates depth estimation errors for better matching. Our proposed MMLF approach significantly improves LT3D performance over prior work, particularly improving rare class performance from 12.8 to 20.0 mAP! Our code and models are available on our project page.
Index Terms:
Long-Tailed Distribution, 3D Detection, Multi-Modal Late-Fusion, Autonomous Vehicles, Open World, LiDAR, RGB.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ 1 Introduction
Few            Medium            Many
Refer to caption Refer to caption
Figure 1: According to the histogram of per-class object counts (on the left), the nuScenes benchmark focuses on the common classes in cyan (e.g., car and barrier) but ignores rare ones in red (e.g., stroller and debris). In fact, the benchmark creates a superclass pedestrian by grouping multiple classes in green, including the common class adult and several rare classes (e.g., child and police-officer); this complicates the analysis of detection performance as pedestrian performance is dominated by adult. Moreover, the ignored superclass pushable-pullable also contains diverse objects such as shopping-cart, dolly, luggage and trash-can as shown in the top row (on the right). We argue that AVs should also detect rare classes as they can affect AV behaviors. Following [1], we report performance for three groups of classes based on their cardinality (split by dotted lines): Many, Medium, and Few.
Three-dimensional (3D) object detection is a key component in many robotics systems such as autonomous vehicles (AVs). To facilitate research in this space, the AV industry has released many large-scale 3D annotated multi-modal datasets [2, 3, 4, 5, 6]. However, these datasets often benchmark on only a few common classes (e.g., car and pedestrian) and ignore rare classes (although annotated by some datasets) like stroller and emergency vehicle (Fig. 1). In the real open world, safe navigation [7, 8] requires AVs to reliably detect rare-class objects such as debris and stroller. This motivates the study of Long-Tailed 3D Detection (LT3D), a problem that requires detecting objects from both common and rare classes.
Status Quo. Among contemporary AV datasets, nuScenes [6] exhaustively annotates objects of various classes crucial to AVs (Fig. 1) and organizes them into a semantic hierarchy (Fig. 4). We are motivated to study LT3D by re-purposing all annotated classes in nuScenes because detecting rare classes is useful for downstream tasks such as motion planning. Importantly, LT3D is not simply solved by retraining state-of-the-art methods on both common and rare classes [9]. For example, CMT [10], a multi-modal transformer-based detector, achieves only 4.8 mAP on rare categories despite achieving 79.9 mAP on common classes (Table I).
Protocol. LT3D requires 3D localization and recognition of objects from each of the common (e.g., adult and car) and rare classes (e.g, child and stroller). Moreover, for safety-critical robots such as AVs, detecting but mis-classifying rare objects (e.g., mis-classifying a child as an adult) is preferable to failing to detect them at all for safe operation. Therefore, we propose a new diagnostic metric to quantify the severity of classification mistakes in LT3D that exploits inter-class relationships w.r.t to the semantic hierarchy when awarding partial credit (Fig. 4). We use both the standard and proposed metrics to evaluate 3D detectors on all classes. Further, since prior work focuses on only a few common classes, they miss opportunities to exploit this semantic hierarchy during training. We propose hierarchical losses to promote feature sharing across both common and rare classes
@@
Figure 2: We extensively explore the simple multi-modal late-fusion (MMLF) framework for LT3D by ensembling RGB and LiDAR uni-modal detectors [9]. We rigorously examine three critical components within this framework (Fig. 3) and propose a simple method that fuses detections produced by a 2D RGB-detector (e.g., DINO [11]) and a 3D LiDAR-detector (e.g., CenterPoint [12]). Our method achieves 51.4 mAP on the nuScenes [6] LT3D benchmark, significantly improving over state-of-the-art detectors by 5.9% (Table I).
@@
Figure 3: We examine three key components in the multi-modal late-fusion (MMLF) of uni-modal RGB and LiDAR detectors from first principles: A. whether to train 2D or 3D RGB detectors, B. whether to match uni-modal detections on the 2D image plane or in the 3D bird’s-eye-view (BEV), and C. how to best fuse matched detections. Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining calibrated scores with Bayesian fusion yields state-of-the-art LT3D performance (Table III and Table IV).
Technical Insights. To address LT3D, we start by retraining state-of-the-art 3D detectors on all classes. Somewhat surprisingly, prior work performs rather poorly on rare classes, e.g., CenterPoint [12] achieves only 0.1 AP on child and 0.1 AP on stroller (Table V). We propose several algorithmic innovations to improve these results. First, we allow feature sharing across common and rare classes by training a single feature trunk, adding in hierarchical losses that ensure features will be useful for all classes (Table IV). Second, noting that LiDAR data is simply too impoverished for even humans to recognize certain tail objects that tend to be small (e.g., strollers), we propose a Multi-Modal Late-Fusion (MMLF) framework (Fig. 2) that fuses detections from a LiDAR-only detector (for precise 3D localization) and an RGB-only detector (for better recognition). We introduce Multi-Modal Filtering (MMF), a simple late-fusion approach that post-processes uni-modal 3D predictions from 3D LiDAR-only (e.g., CenterPoint [12]) and 3D RGB-only (e.g., FCOS3D [13]) detectors, filtering away detections that are inconsistent across modalities. This significantly improves LT3D performance by 3.2 mAP on average (Table II). Next, we delve into this MMLF framework and study three critical design choices (Fig. 3): (A) whether to train a 2D or a 3D monocular RGB detector for late-fusion, (B) whether to match detections in the 2D image plane or in 3D, and (C) how to optimally fuse detections. Our exploration reveals that using 2D RGB detectors, matching on the 2D image plane, and combining score-calibrated predictions with Bayesian fusion yields state-of-the-art LT3D performance, significantly outperforming end-to-end trained multi-modal 3D detectors (Table I).
Contributions. We present three major contributions. First, we formulate the problem of LT3D, emphasizing the detection of both common and rare classes in safety-critical applications like AVs. Second, we design LT3D’s benchmarking protocol and develop a diagnostic metric that awards partial credit depending on the severity of misclassifications (e.g., misclassifying child-vs-adult is less problematic than misclassifying child-vs-car). Third, we propose several architecture-agnostic approaches to address LT3D, including a simple multi-modal late-fusion (MMLF) strategy that generalizes across different RGB and LiDAR architectures. We conduct extensive experiments to ablate our design choices and demonstrate that our simple MMLF approach achieves state-of-the-art results on the nuScenes and Argoverse 2 LT3D benchmarks.
@@ 2 Related Work
3D Object Detection for AVs. Contemporary approaches to 3D object detection can be broadly classified as LiDAR-only, RGB-only, and sensor-fusion methods. Recent work in 3D detection is heavily inspired by prior work in 2D detection [14, 15, 16]. LiDAR-based detectors like PointPillars [17], CBGS [18], and PVRCNN++ [19] adopt an SSD-like architecture [15] that regresses amodal bounding boxes from a bird’s-eye-view (BEV) feature map. More recently, CenterPoint [12] adopts a center-regression loss that is inspired by CenterNet [14]. Despite significant progress, LiDAR-based detectors often produce many false positives because it is difficult to distinguish foreground objects from the background given sparse LiDAR returns. Monocular RGB-based methods have gained popularity in recent years due to increased interest in camera-only perception. FCOS3D [13] extends FCOS [20] by additionally regressing the size, depth, and rotation for each object. More recently, methods such as BEVDet and BEVFormer [21, 22, 23] construct a BEV feature-map by estimating the per-pixel depth of each image feature [24]. PolarFormer [25] introduces a polar-coordinate transformation that improves near-field detection. Importantly, many of these state-of-the-art 3D RGB detectors are commonly pre-trained on large external datasets like DDAD [26]. Monocular RGB detectors accurately classify objects but struggle to estimate depth, particularly for far-field detections [27]. Despite recent advances in LiDAR and RGB 3D detectors, we find that multi-modal fusion is essential for LT3D (detailed next). Importantly, using both RGB (for better recognition) and LiDAR (for better 3D localization) helps detect rare classes. We study the multi-modal late-fusion framework described in Fig. 2 to determine how to effectively fuse RGB and LiDAR uni-modal detectors for LT3D.
@@
Multi-modal 3D Detection. Conventional wisdom suggests that fusing multi-modal cues, particularly using LiDAR and RGB, can improve 3D detection. Intuitively, LiDAR faithfully measures the 3D world (although it has notoriously sparse point returns), and RGB has high-resolution that captures semantic features for recognition (but lacks 3D information). Multi-modal fusion for 3D detection is an active field of exploration. Popular approaches can be categorized as input-fusion, feature-fusion, and late-fusion. Input-fusion methods typically augment LiDAR points using image-level features. For example, PointPainting [28] projects LiDAR points onto the output mask of a semantic segmentation model and appends corresponding class scores to each point. MVP [29] densifies regions of LiDAR sweeps that correspond with objects in semantic segmentation masks. In contrast, Frustum PointNets [30] leverage 2D RGB detections to localize objects within the box frustum using PointNets [31].
@@
Recent works show that feature-fusion can be more effective than input-fusion. PointFusion [32] fuses global image and point-cloud features prior to detection and MSMDFusion [33] fuses LiDAR and RGB features at multiple scales. TransFusion [34] and BEVFusion [35] fuse features in the BEV space using multi-headed attention. Despite the success of transformers for detecting common objects, [9] finds that TransFusion struggles to detect rare classes, and posits that the transformer architecture, as adopted in TransFusion and BEVFusion, suffers from limited training data (particularly for classes in the long tail). For transformers to work well in practice, they should be trained on diverse, large-scale datasets [36, 37]. It is also worth noting that end-to-end trained multi-modal detectors require paired multi-modal data for training, increasing the cost of data collection and modality alignment. In this work, we opt to study a multi-modal late-fusion strategy, a framework that ensembles uni-modal detectors, which do not require aligned RGB-LiDAR paired training data. CLOCs [38] is a late-fusion method that learns a separate network to fuse RGB and LiDAR detections, showing promising results for 3D detection. Importantly, prior late-fusion methods like CLOCS [38] only perform late-fusion on matched predictions with semantic agreement and do not fix misclassifications. We find that handling such misclassifications is critical for improving rare class performance. We delve into this simple multi-modal late-fusion framework, study three crucial design choices, and present a method that significantly outperforms the state-of-the-art for LT3D.
Long-Tailed Perception. AV datasets follow a long-tailed class distribution: a few classes like car and pedestrian are dominant, while others like stroller and debris are rarely seen. However, this problem is not unique to the AV domain [39] – Long-Tailed Perception (LTP) is a long-standing problem in the literature [1] and has been widely studied through the lens of image classification, aiming for high accuracy averaged across imbalanced classes [1, 40, 41]. Existing LTP methods propose reweighting losses [42, 43, 44, 45, 46, 47], rebalancing data sampling [48, 49, 50], balancing gradients computed from imbalanced classes [51], and balancing network weights [41]. Others study LTP through the lens of 2D object detection with RGB images [52]. Compared to 2D image-based recognition, long-tailed 3D detection has unique opportunities and challenges because sensors such as LiDAR directly provide geometric and ego-motion cues that are difficult to extract from 2D images. Further, 2D detectors must detect objects of different scales due to perspective image projection, dramatically increasing the complexity of the output space (e.g., requiring more anchor boxes). In contrast, 3D objects do not exhibit as much scale variation, but far-away objects tend to have sparse LiDAR returns [27, 53], imposing different challenges. Finally, 3D detectors often use class-aware heads (i.e. each class has its own binary classifier) while 2D long-tail recognition approaches typically use shared softmax heads.
Recently, CBGS [18] explicitly addresses rare-class 3D detection by up-sampling LiDAR-sweeps with instances of rare classes, and pasting instances of rare objects copied from different scenes. Although this works well for improving detection of infrequently-seen classes (e.g. classes with medium number of examples like bicycle and construction vehicle), it does not provide significant improvement for classes with only a few examples like debris and stroller [9]. Additionally, rare classes, such as child and stroller, are typically small in size and have a limited number of LiDAR returns. As a result, LiDAR-only detectors struggle to accurately recognize these rare classes. In this work, we address LT3D by multi-modal late-fusion, which ensembles RGB and LiDAR uni-modal detectors.
3 LT3D Benchmarking Protocol
Conceptually, LT3D extends the traditional 3D detection problem, which focuses on identifying objects from $K$ common classes, by further requiring detection of $N$ rare classes. As LT3D emphasizes detection performance on all classes, we report metrics for three groups of classes based on their cardinality (Fig. 1-left): many ($>$50k instance/class), medium (5$\sim$50k instance/class), and few ($<$5k instance/class). To better analyze LT3D performance, we present two metrics below.
Mean average precision (mAP) is a well-established metric for object detection [54, 2, 55]. For 3D detection on LiDAR sweeps, a true positive is defined as a detection that has a center distance within a distance threshold to a ground-truth annotation [6]. mAP computes the mean of AP over classes, where per-class AP is the averaged area under the precision-recall curves with distance thresholds of [0.5, 1, 2, 4] meters.
Hierarchical Mean Average Precision (mAPH). For safety-critical applications, it is more desirable to localize-but-misclassify an object than miss-detect this object, e.g., detecting but misclassifying a child as adult is better than not detecting this child. Therefore, we present hierarchical mean average precision (mAPH), which considers such semantic relationships across classes to award partial credit. We use APH as a supplementary diagnostic tool to analyze how LT3D detectors make mistakes. To encode inter-class relationships, we leverage the semantic hierarchy (Fig. 4) such as the one defined by nuScenes [6]. We derive partial credit as a function of semantic similarity using the least common ancestor (LCA) distance metric. LCA has been proposed for image classification [56, 57] but not for object detection. Extending this metric to object detection is challenging as we must jointly evaluate semantic and spatial overlap. For clarity, we describe how to compute APH for a class $C$.
LCA=0: Consider the predictions and ground-truth boxes for $C$. Label the set of predictions that overlap with ground-truth boxes for $C$ as true positives. Other predictions are false positives. This is identical to the standard mAP metric.
LCA=1: Consider the predictions for $C$, and ground-truth boxes for $C$ and its all sibling classes with LCA distances to $C$ of 1. Label the predictions that overlap a ground-truth box of $C$ as a true positive. Label the predictions that overlap sibling classes as ignored [55]. All other predictions for $C$ are false positives.
LCA=2: Consider the predictions for $C$ and ground-truth boxes for $C$ and all its sibling classes with LCA distances to $C$ no greater than $2$. For nuScenes, this includes all classes. Label the set of predictions that overlap ground-truth boxes for $C$ as true positives. Label the set of predictions that overlap other classes as ignored. All other predictions for $C$ are false positives.
4 LT3D Methods: The Devil Is In The Details
To address LT3D, we first retrain state-of-the-art (SOTA) detectors on all long-tailed classes. We consider the following detectors:
  • LiDAR-based 3D Detectors: CenterPoint [12], PointPillars [17], TransFusion-L [34], BEVFusion-L [35], CMT-L [10];
  • RGB-based 3D Detectors: FCOS3D [13], PolarFormer [25], BEVFormer [23];
  • RGB-based 2D Detectors: YOLOV7 [58], and DINO [11];
  • End-to-End Multi-modal 3D Detectors: TransFusion [34], BEVFusion [35], CMT [10], and DeepInteraction [59].
In Section 4.1 and 4.2, we introduce two modifications that consistently improve LT3D performance, and describe our multi-modal late-fusion framework.
Figure 4: The nuScenes dataset defines a semantic hierarchy for all annotated classes (Fig. 1). We highlight common classes in white and rare classes in gold. The standard nuScenes benchmark makes two choices for dealing with rare classes: (1) ignore them (e.g., stroller and pushable-pullable), or (2) group them into coarse-grained classes (e.g., adult, child, construction-worker, police-officer are grouped as pedestrian). Since the pedestrian class is dominated by adult (Fig. 1), the standard benchmarking protocol masks the challenge of detecting rare classes like child and police-officer. We explicitly incorporate this semantic hierarchy to evaluate LT3D performance via the proposed mAPH metric described in Sec. 3.
4.1 Grouping-Free Detector Head
Extending existing 3D detectors to train with more classes is surprisingly challenging. Many contemporary networks use a multi-head architecture that groups classes of similar size and shape to facilitate efficient feature sharing [18, 12]. For example, CenterPoint [12] groups pedestrian and traffic-cone since these objects are both tall and skinny. However, this multi-head grouping design may not work for super-classes like pushable-pullable and debris that contain diverse objects of different sizes and shapes. Moreover, in the multi-head design, each head is a group-specific detector that consists of several layers with lots of parameters. This makes multi-head grouping difficult to scale for a large number of classes. To address these issues, we treat each class as its own group to avoid hand-crafted grouping heuristics and design a group-free strategy, in which each class has only one linear layer as its detector (Fig. 7) and all classes share a single detector head. This design significantly reduces the number of parameters and allows learning the shared feature backbone collaboratively with all classes, effectively mitigating overfitting to rare classes. Adding a new class is as simple as adding a single linear layer to the detector head. Importantly, we demonstrate that our grouping-free detector head outperforms conventional grouping-based methods in the supplement.
4.2 Training with a Semantic Hierarchy
The nuScenes dataset defines a semantic hierarchy (Fig. 4) for all classes, grouping semantically similar classes under coarse-grained categories [6]. We leverage this hierarchy during training. Specifically, we train detectors to predict three labels for each object: its fine-grained class (e.g., child), its coarse class (e.g., pedestrian), and the root class object. We adopt a grouping-free detector head (Fig. 7) that outputs separate “multitask” heatmaps for each class, and use a per-class sigmoid focal loss [60] rather than multi-class cross-entropy loss to avoid normalizing class probabilities across fine-grained and super-classes. It is worth noting that our approach does not explicitly enforce a tree-like hierarchy, and can be applied to more complex label relationships [61]. Crucially, adding a vehicle heatmap does not directly interfere with the car heatmap. However, this might produce repeated detections on the same test object (e.g. a single ground-truth car may be detected as a car, vehicle, and object). We address that by simply ignoring coarse detections at test time. We explore alternatives in the supplement and conclude that they achieve similar LT3D performance. Perhaps surprisingly, this training method improves detection performance not only for rare classes, but also for common classes (Table IV).
4.3 Multimodal Filtering for Detection Fusion
Rare class instances are often small and can be challenging to recognize from sparse (LiDAR) geometry alone – even humans struggle to find strollers in LiDAR point clouds. This suggests that one can leverage multi-modal cues to improve LT3D. First, we find that, although LiDAR-based detectors are widely adopted for 3D detection, they produce many high-scoring false positives (FPs) for rare classes due to misclassification. We focus on removing such FPs. To this end, we use an RGB-based 3D detector (e.g., FCOS3D [13]) to filter out high-scoring false-positive LiDAR detections by leveraging two insights: (1) LiDAR-based 3D-detectors can achieve high recall and precise 3D localizations for TPs, and (2) RGB-based 3D-detections are accurate w.r.t recognition although their 3D localization is poor. Fig. 5 demonstrates our Multi-Modal Filtering (MMF) strategy. For each RGB detection, we keep LiDAR detections within a radius of $m$ meters and remove all the other LiDAR detections. We denote this method as MMF($D_{L}$, $D_{R}$), where DL and $D_{R}$ are any 3D LiDAR detector and 3D RGB detector, respectively. Table II demonstrates that MMF greatly improves LT3D.
LiDAR Detections       RGB Detections          Filtered Results
Figure 5: Multi-Modal Filtering (MMF) effectively removes high-scoring false-positive LiDAR detections. The green boxes are ground-truth strollers, while the blue boxes are stroller detections from the LiDAR-based detector CenterPoint [12] (left) and RGB-based detector FCOS3D [13] (mid). The final filtered result removes LiDAR detections not within $m$ meters of any RGB detection, shown in the red region, and keeps all other LiDAR detections, shown in the white region (right).
Vehicle             Pedestrian           Movable
Refer to caption Refer to caption
Figure 6: We analyze our multi-modal filtering (MMF) approach. Rare classes are most often confused by the dominant class (in blue) in each superclass: (left) Vehicle is dominated by car, (mid) Pedestrian is dominated by adult, and (right) Movable is dominated by barrier. We find that class confusions are reasonable. Car is often mistaken for truck. Similarly, truck, construction-vehicle and emergency-vehicle are most often mistaken for car. Bicycle and motorcycle are sometimes misclassified as car, presumably because they are sometimes spatially close to cars. Adults have similar appearance to police-officer and construction-worker, and they are often co-localized with child and stroller.
Limitations of Multi-Modal Filtering. Despite the effectiveness of multi-modal filtering, this approach is sensitive to the classification accuracy of FCOS3D when matching LiDAR and RGB-based detections, leading to many correctly localized but misclassified detections. We use a confusion matrix to further analyze the misclassifications within each superclass, as shown in Fig. 6. We explain how to compute a confusion matrix for the detection task: For each superclass, we construct a confusion matrix, in which the entry $(i,j)$ indicates the misclassification rate of class-$i$ objects as class-$j$. Specifically, given a class $i$, we find its predictions that match ground-truth boxes within a 2m center-distance of class-$i$ and all its sibling classes (LCA=1, within the corresponding superclass); we ignore all unmatched detections. This allows us to count the mis-classifications of class-$i$ objects into class-$j$. We normalize these counts to produces misclassification rates. We find that rare classes are most often confused by the dominant class in each superclass: Vehicle is dominated by car, Pedestrian is dominated by adult, and Movable is dominated by barrier. We posit that this can be addressed with more accurate RGB-based classifiers, motivating our study of multi-modal late-fusion (MMLF) from first principles.
Figure 7: We leverage the semantic hierarchy defined in the nuScenes dataset (Fig. 4) to train LT3D detectors by predicting class labels at multiple levels of the hierarchy for an object: its fine-grained label (e.g., child), its coarse class (e.g., pedestrian), and the root-level class object. This means that the final vocabulary of classes is no longer mutually exclusive, so we use a sigmoid focal loss that learns separate spatial heatmaps for each class.
4.4 Delving into Multi-Modal Late-Fusion for LT3D
Motivated by the effectiveness of our frustratingly simple multi-modal filtering (MMF) approach, we delve into this multi-modal late-fusion (MMLF) strategy. We explore three questions from first principles (Fig. 3): (1) how do we effectively incorporate RGB information, (2) how do we match RGB and LiDAR detections, and (3) how do we fuse them. We denote our final approach as MMLF($D_{L}$, $D_{R}$) in Table I, where DL is any 3D LiDAR detector and $D_{R}$ is any 2D RGB detector.
4.4.1 How Do We Incorporate RGB Information?